OpenStudy (anonymous):

trying to solve this simple quadratic equation, can any one help find the correct factors to find the solution. x^2 + x + 6 = 0

OpenStudy (amistre64):

id say it has something to do with 3 and 2

OpenStudy (amistre64):

what factors of the last number add to get the middle?

OpenStudy (anonymous):

Ive tried each possible solution but I cannot come up with one because I cannot place A - sign and come up with a pos 6 at the end

OpenStudy (amistre64):

x^2 + x + 6 = 0 this is the correct equation right?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

if so we factor it out to be (x+3)(x+2) = 0; not the rule of math is; anything times zero = zero which implies that when x+3 = 0 or x+2 = 0 we have good answers

OpenStudy (amistre64):

so when x=-3 or x=-2 we have y=0

OpenStudy (amistre64):

not the rule meant now the rule .... cant type

OpenStudy (anonymous):

so ....x = 0

OpenStudy (amistre64):

i cant imagine where you come up with that from what i wrote .....

OpenStudy (anonymous):

from confusion

OpenStudy (amistre64):

so when x=-3 or x=-2 we have y=0 x=0 is nowhere to be read in that. the answers are: x = -3 and x = -2 but answers dont really help you understand the problem itself

OpenStudy (amistre64):

quadratics are not intuitive simply becasue they are curved; and we are used to dealing with straight lines that only have 1 solution

OpenStudy (amistre64):

quadratics have at most 2 solutions

OpenStudy (amistre64):



OpenStudy (amistre64):

 sometimes they even have NO real solutions